Estimate the number of electrons in 100 g of water. How much is the total negative charge on these electrons?

Grade:upto college level

2 Answers

Navjyot Kalra
9 years ago
Sol. No. of electrons of 100g water = 10 * 100/18 = 55.5 Nos total charge = 55.5 No. of electrons in 18 g of H base 2O = 6.023 * 10^23 * 10 = 6.023 * 10^24 No. of electrons in 100 g of H base 2O = 6.023 * 24^24 * 100/18 = 0.334 * 10^26 = 3.334 * 10^25 Total charge = 3.34 * 10^25 * 1.6 * 10^-19 = 5.34 * 10^6 c
Arun Kumar IIT Delhi
9 years ago
Hello
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty